# Learnability is a Compact Property

### v1

@@ 2.3 Transductive Learning
where $A(S_{-i},h)$ denotes the output of $A$ on the sample $(x_{j},h(x_{j}))_{x_{j}\in S_{-i}}$. Informally, the setting of transductive learning can be thought of as a fine-grained analogue of PAC learning in which learners are judged on the basis of their worst-case performance on a sample rather than over a distribution.
@@
The transductive error rate of a learner $A$ for $\mathcal{H}$ is the function $\xi_{A,\mathcal{H}}:\mathbb{N}\to\mathbb{R}$ defined by $\xi_{A,\mathcal{H}}(n)=\sup_{S\in\mathcal{X}^{n},\;h\in\mathcal{H}}L_{S,h}^{\operatorname{Trans}}(A)$. The transductive sample complexity of a learner $A$ for $\mathcal{H}$ is the function
$m_{\operatorname{Trans},A}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{A,\mathcal{H}}(m^{\prime})<\epsilon,\;\forall m^{\prime}\geq m\}.$
@@
The transductive error rate of a class $\mathcal{H}$ is the minimal error rate attained by any of its learners, i.e., $\xi_{\mathcal{H}}(n)=\inf_{A}\xi_{A,\mathcal{H}}(n)$. The transductive sample complexity $m_{\operatorname{Trans},\mathcal{H}}:(0,1)\to\mathbb{N}$ of $\mathcal{H}$ is the function mapping $\epsilon$ to the minimal $m$ for which $\xi_{\mathcal{H}}(m^{\prime})<\epsilon$ for all $m^{\prime}\geq m$. That is,
$m_{\operatorname{Trans},\mathcal{H}}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{\mathcal{H}}(m^{\prime})<\epsilon,\;\forall m^{\prime}\geq m\}.$
The result which justifies our emphasis on transductive learnability — and furthermore on transductive error rates and sample complexity — is the equivalence between transductive and PAC sample complexities, up to modest factors.
Lemma 2.6 (Asilis et al. [7, Proposition 3.6]).
Let $\mathcal{X}$ be a domain, $\mathcal{Y}$ a label set, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Fix a loss function taking values in $[0,1]$. Then the following inequality holds for all $\epsilon\in(0,1)$ and an absolute constant $c>0$:
$m_{\operatorname{Trans},\mathcal{H}}(c\cdot\epsilon)\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\epsilon)\leq m_{\operatorname{Trans},\mathcal{H}}(\epsilon/c)\cdot c\cdot\log(1/\epsilon).$
See Asilis et al. [7, Section 3] for further detail on the equivalence between minimizing transductive error, PAC error (i.e., providing high-probability guarantees), and expected error.
@@ 3 Compactness of Learning
We present the central result of the paper in this section: learnability, and moreover the sample complexity of learning, are compact properties of a hypothesis class. In Section 3.1 we study compactness of supervised learning over proper loss functions, and demonstrate a strong compactness result: a class $\mathcal{H}$ is learnable with transductive sample complexity $m$ if and only if all its finite projections are learnable with the same complexity. In Section 3.2 we examine the case of supervised learning over improper loss functions and prove a negative result: the previous compactness result no longer holds in this more general setting. Nevertheless, we are able to demonstrate compactness for the special case of the (improper) 0-1 loss function, i.e., multiclass classification over arbitrary, possibly infinite label sets. Notably, this recovers M. Hall’s classic matching theorem for infinite graphs [24] as a corollary to our central result.
In Sections 3.3 and 3.4 we depart from learning in the realizable case and examine analogues of our results for agnostic learning (Section 3.3) and distribution-family learning (Section 3.4).
3.1 Proper Loss Functions
We first consider the case of loss functions $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ which endow the label set $\mathcal{Y}$ with the structure of a proper metric space.
@@ 3.1 Proper Loss Functions
We remark that proper spaces are sometimes referred to as Heine-Borel spaces, and that their examples include $\mathbb{R}^{n}$ endowed with any norm, all closed subsets of $\mathbb{R}^{n}$, and all compact metric spaces (e.g., all finite sets endowed with arbitrary metrics). Further discussion of proper metric spaces is provided in Appendix A. The central technical result of this subsection is a compactness property concerning assignments of variables to metric spaces that maintain a family of functions below a target value $\epsilon$. As we will see shortly, compactness of supervised learning with proper label spaces arises as a consequence of this property.
@@
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. Before arguing the reverse direction, some terminology: a partial assignment of variables is an assignment of variables for a subset of $L$. A partial assignment is said to be completable with respect to $R^{\prime}\subseteq R$ if its unassigned variables can all be assigned so that the functions $r^{\prime}\in R^{\prime}$ are kept below $\epsilon$. A partial assignment is finitely completable if it is completable with respect to all finite subsets of $R$. Notably, this is a pointwise condition, not a uniform one. The completions are permitted to vary across the finite subsets of $R$.
@@
To see why, let $|R^{\prime}|=k$ and let $\ell_{1},\ldots,\ell_{m}$ be the variables in $L$ upon which the functions in $R^{\prime}$ depend. Suppose without loss of generality that $\ell=\ell_{1}$ and that nodes $\ell_{i+1},\ldots,\ell_{m}$ have already been assigned. Now consider the function $f_{R^{\prime}}:\prod_{j=1}^{i}M_{\ell_{j}}\to\mathbb{R}^{k}$ mapping assignments of the $\ell_{1},\ldots,\ell_{i}$ to the outputs they induce on the functions in $R^{\prime}$. (Notably, this includes the assignments already made for $\ell_{i+1},\ldots,\ell_{m}$.) As the functions in $R^{\prime}$ are proper, including when some of their inputs are fixed, the function $f_{R^{\prime}}$ is proper as well.
Thus $f_{R^{\prime}}^{-1}([0,\epsilon]^{k})$ is compact, along with its its projection onto its first coordinate. That set is precisely $N_{R^{\prime}}(\ell)$, demonstrating our intermediate claim. We thus have a family of compact, non-empty sets
=$\mathcal{I}=\left\{N_{R^{\prime}}(\ell):R^{\prime}\subseteq R,|R^{\prime}|\leq\infty\right\}.$
Note that finite intersections of elements of $\mathcal{I}$ are non-empty, as
$\bigcap_{i=1}^{j}N_{R_{i}}(\ell)\supseteq N_{\bigcup_{i=1}^{j}R_{i}}(\ell)\neq\emptyset.$
@@
Thus, invoking Zorn’s lemma, $\mathcal{P}$ has a maximal element $\phi_{\max}$. By Lemma 3.4, it must be that $\phi_{\max}$ does not leave a single variable unassigned, otherwise it could be augmented with an additional assignment. There thus exists a total assignment that is finitely completable. As it has no free variables, it must indeed be maintaing all functions in $R$ below $\epsilon$. The claim follows. ∎
We remark that Theorem 3.3 can be seen to fail miserably if topological requirements are removed from the functions in $R$. For instance, let there be a single variable $L=\{\ell\}$ which takes values in $\mathbb{N}$ and countably many functions $R=\{r_{i}\}_{i\in\mathbb{N}}$ with $r_{i}(\ell)=[\ell=i]$. Then any finite $R^{\prime}\subseteq R$ can be made to evaluate to 0 by setting $\ell$ to a sufficiently large natural number, but any assignment for $\ell$ guarantees that some function in $R$ evaluate to 1. In fact, if one permits the functions in $R$ to be improper, a counterexample can furthermore be found in which each function $\ell\in L$ influences only finitely many functions in $R$ (though this is not a requirement of Theorem 3.3). Namely, let $L=\{\ell_{i}\}_{i\in\mathbb{N}}$ where each variable resides in $\mathbb{N}$, and let $R=\{r_{i}\}_{i\in\mathbb{N}}$ with $r_{i}(x_{i},x_{i+1})=[x_{i}\leq x_{i+1}]$. Then each variable influences at most two functions and each finite collection $\{r_{1},\ldots,r_{n}\}$ can be maintained at 0 by assigning $x_{1}>x_{2}>\ldots>x_{n}$, but it is not possible to have $x_{i}>x_{i+1}$ for all $i$. We leave open the mathematically intriguing question of whether the requirement that each $r\in R$ depend upon finitely many variables in $L$ can be removed, provided that all functions in $R$ are proper.
@@
A corollary to Theorem 3.3 is that the same claim holds when the target values $\epsilon$ vary over the functions $r\in R$, as translations and scalings of proper functions $Z\to\mathbb{R}$ are proper. (In particular, the compact subsets of $\mathbb{R}$ are closed under translation and scaling.)
@@
Let $\mathcal{X}$ be an arbitrary domain, $\mathcal{Y}$ a label set with the structure of a proper metric space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Then the following conditions are equivalent:
@@
    $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the realizable case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. For the reverse direction, fix an $\epsilon\in(0,1)$ and set $n=m(\epsilon)$. Then fix a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$. It suffices to demonstrate that a transductive learner for $\mathcal{H}$ on instances of the form $\{(S,h)\}_{h\in\mathcal{H}}$ can be designed which attains error $\leq\epsilon$.
We will capture the task of transductively learning $\mathcal{H}$ on such instances by way of a certain collection $R$ of functions and $L$ of variables. Each variable $\ell\in L$ will be permitted to take values in $\mathcal{Y}$, while each function $r\in R$ depends upon exactly $n$ variables in $L$ and outputs values in $\mathbb{R}_{\geq 0}$. More precisely, let $R=\mathcal{H}|_{S}$ and $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$. These serve merely as representations for the functions in $R$ and variables in $L$, not their true definitions (which will be established shortly). Note now that by supressing the unlabeled datapoints of $S$, we can equivalently represent elements of $R$ as sequences in $\mathcal{Y}^{n}$ and elements of $L$ as sequences in $(\mathcal{Y}\cup\{?\})^{n}$. In particular, each element of $L$ is precisely an element of $R$ which had exactly one entry replaced with a ``?"``?"``?"` ` ? ".
Now, to model transductive learning, fix a node $r\in R$ represented by $(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$. Then we will define $r$ to be a function depending upon the variables $\ell_{1},\ldots,\ell_{n}\in L$, where
$\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n}).$
Given assignments for each of the variables $\ell_{1},\ldots,\ell_{n}$ as values in $\mathcal{Y}$ — semantically, completions of their “?” entries — the node $r$ then outputs the value $\frac{1}{n}\cdot\sum_{i=1}^{n}d_{\mathcal{Y}}(y_{i},\ell_{i})$. The two crucial observations are as follows: an assignment of each $\ell\in L$ corresponds precisely to the action of a learner responding to a query at test time, and the output of node $r$ equals the error of a learner when $r$ is the ground truth.
@@
3.2 Improper Loss Functions
It is natural to ask whether the requirement that $\mathcal{Y}$ be a proper metric space is essential to Theorem 3.6 or merely an artifact of the proof. We now demonstrate the former: for arbitrary metric losses, the error rate of learning $\mathcal{H}$ can exceed that of all its finite projections by a factor of 2.
@@ 3.2 Improper Loss Functions
There exists a metric space $\mathcal{Y}$ and hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ such that:
  • For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\xi_{\mathcal{H}^{\prime}}\leq\xi$ for a fixed $\xi:\mathbb{N}\to\mathbb{R}_{\geq 0}$.
  • There exists $n\in\mathbb{N}$ such that $\xi_{\mathcal{H}}(n)\geq 2\cdot\xi(n)$.
Let us describe the main idea of Theorem 3.7, whose proof is deferred to Appendix B.1. The crucial step lies in the creation of the label space $\mathcal{Y}=R\cup S$, where $R$ is an infinite set whose points are all distance 2 apart, and $S$ is an infinite set whose elements are indexed by the finite subsets of $R$, e.g., as in $s_{R^{\prime}}\in S$ for finite $R^{\prime}\subseteq R$. For all such $R^{\prime}$, define $s_{R^{\prime}}$ to be distance 1 from the elements of $R^{\prime}$, distance 2 from the other elements of $R$, and distance 1 from all other points in $S$. Then $\mathcal{Y}$ indeed forms a metric space, and it is straightforward to see that, for instance, $\mathcal{H}=\mathcal{Y}^{[1]}$ is more difficult to learn than its finite projections (equivalently, finite subsets).
We now prove a matching upper bound to Theorem 3.7, demonstrating that a factor of 2 is the greatest possible gap between the error rate of $\mathcal{H}$ and its projections when $\mathcal{Y}$ is a metric space.

### v2

@@ 2.2 Transductive Learning
where $A(S_{-i},h)$ denotes the output of $A$ on the sample $(x_{j},h(x_{j}))_{x_{j}\in S_{-i}}$.
@@
The transductive error rate of a learner $A$ with respect to a class $\mathcal{H}$ is the function $\xi_{A,\mathcal{H}}(n)=\sup_{S\in\mathcal{X}^{n},\;h\in\mathcal{H}}L_{S,h}^{\operatorname{Trans}}(A)$. The transductive sample complexity of $A$ is the function
$m_{\operatorname{Trans},A}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{A,\mathcal{H}}(m^{\prime})\leq\epsilon,\;\forall m^{\prime}\geq m\}.$
@@
The transductive error rate of a class $\mathcal{H}$ is the minimal error rate attained by any of its learners, i.e., $\xi_{\mathcal{H}}(n)=\inf_{A}\xi_{A,\mathcal{H}}(n)$. The transductive sample complexity $m_{\operatorname{Trans},\mathcal{H}}:\mathbb{R}_{>0}\to\mathbb{N}$ of $\mathcal{H}$ is the function mapping $\epsilon$ to the minimal $m$ for which $\xi_{\mathcal{H}}(m^{\prime})\leq\epsilon$ for all $m^{\prime}\geq m$. That is,
$m_{\operatorname{Trans},\mathcal{H}}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{\mathcal{H}}(m^{\prime})\leq\epsilon,\;\forall m^{\prime}\geq m\}.$
Informally, agnostic transductive learning is the analogue in which the adversary is permitted to label the data in $S$ arbitrarily, and in which the learner need only compete with the best hypothesis in $\mathcal{H}$. We defer the formal definition to Section 3.3.
@@ 3 Compactness of Learning
We present the central result of the paper in this section: the transductive sample complexity of learning is a compact property of a hypothesis class. In Section 3.1 we study compactness of realizable supervised learning over proper loss functions, and demonstrate a strong compactness result: a class $\mathcal{H}$ is learnable with transductive sample complexity $m$ if and only if all its finite projections are learnable with the same complexity. In Section 3.2 we examine the case of realizable supervised learning over improper loss functions and prove a negative result: the previous compactness result no longer holds in this more general setting. Nevertheless, we demonstrate an approximate form of compactness, up to a factor of 2, for (improper) metric losses. Moreover, we show exact compactness for the special case of the (improper) 0-1 loss function, i.e., multiclass classification over arbitrary, possibly infinite label sets. Notably, this recovers M. Hall’s classic matching theorem for infinite graphs (Hall Jr, 1948) as a corollary to our central result. In Section 3.3 we examine analogues of our results for agnostic learning, and in Section 3.4 we transfer our results to the PAC model via standard equivalences, obtaining approximate compactness of sample complexities. Finally, in Section 3.5 we extend our results to certain forms of distribution-family learning.
3.1 Realizable Learning With Proper Loss Functions
We first consider the case of loss functions $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ defined on a proper metric space $\mathcal{Y}$.
@@ 3.1 Realizable Learning With Proper Loss Functions
We remark that proper spaces are sometimes referred to as Heine-Borel spaces, and that their examples include $\mathbb{R}^{d}$ endowed with any norm, all closed subsets of $\mathbb{R}^{d}$ (under the same norms), and all finite sets endowed with arbitrary metrics. Further discussion of proper metric spaces is provided in Appendix A. The central technical result of this subsection is a compactness property concerning assignments of variables to metric spaces that maintain a family of functions below a target value $\epsilon$.
@@
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. Before arguing the reverse direction, some terminology: a partial assignment of variables is an assignment of variables for a subset of $L$. A partial assignment is said to be completable with respect to $R^{\prime}\subseteq R$ if its unassigned variables can all be assigned so that the functions $r^{\prime}\in R^{\prime}$ are kept below $\epsilon$. A partial assignment is finitely completable if it is completable with respect to all finite subsets of $R$. Notably, this is a pointwise condition, not a uniform one; the completions are permitted to vary across $R$’s subsets.
@@
To see why, let $|R^{\prime}|=k$ and let $\ell_{1},\ldots,\ell_{m}$ be the variables in $L$ upon which the functions in $R^{\prime}$ depend. Suppose without loss of generality that $\ell=\ell_{1}$ and that nodes $\ell_{i+1},\ldots,\ell_{m}$ have already been assigned. Consider the function $f_{R^{\prime}}:\prod_{j=1}^{i}M_{\ell_{j}}\to\mathbb{R}^{k}$ mapping assignments of the $\ell_{1},\ldots,\ell_{i}$ to the outputs they induce on the functions in $R^{\prime}$. (Notably, this includes the assignments already made for $\ell_{i+1},\ldots,\ell_{m}$.) As the functions in $R^{\prime}$ are proper, including when fixing some of their inputs, $f_{R^{\prime}}$ is as well.
Thus $f_{R^{\prime}}^{-1}([0,\epsilon]^{k})$ is compact, as is its projection onto its first coordinate. That set is precisely $N_{R^{\prime}}(\ell)$, demonstrating our intermediate claim. We thus have a family of compact, non-empty sets =$\mathcal{I}=\left\{N_{R^{\prime}}(\ell):R^{\prime}\subseteq R,|R^{\prime}|\leq\infty\right\}$. Note that finite intersections of elements of $\mathcal{I}$ are non-empty, as $\bigcap_{i=1}^{j}N_{R_{i}}(\ell)\supseteq N_{\bigcup_{i=1}^{j}R_{i}}(\ell)\neq\emptyset$.
@@
Thus, invoking Zorn’s lemma, $\mathcal{P}$ has a maximal element $\phi_{\max}$. By Lemma 3.4, it must be that $\phi_{\max}$ does not leave a single variable unassigned, otherwise it could be augmented with an additional assignment. There thus exists a total assignment that is finitely completable. As it has no free variables, it must indeed be maintaining all functions in $R$ below $\epsilon$. The claim follows. ∎
@@
A corollary to Theorem 3.3 is that the same claim holds when the target values $\epsilon$ vary over the functions $r\in R$, as translations and scalings of proper functions $Z\to\mathbb{R}$ are proper.
@@
Let $\mathcal{X}$ be an arbitrary domain, $\mathcal{Y}$ a label set, and $d$ a loss function such that $(\mathcal{Y},d)$ is a proper metric space. Then the following are equivalent for any $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and $m\colon\mathbb{R}_{>0}\to\mathbb{N}$:
@@
    $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the realizable case with transductive sample function $m$.
@@
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. For the reverse direction, fix an $\epsilon>0$ and set $n=m(\epsilon)$. Then fix a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$. It suffices to demonstrate that a transductive learner for $\mathcal{H}$ on instances of the form $\{(S,h)\}_{h\in\mathcal{H}}$ can be designed which attains error $\leq\epsilon$.
We will capture the task of transductively learning $\mathcal{H}$ on such instances by way of a certain collection $R$ of functions and $L$ of variables. Each variable $\ell\in L$ will be permitted to take values in $\mathcal{Y}$, while each function $r\in R$ depends upon exactly $n$ variables in $L$ and outputs values in $\mathbb{R}_{\geq 0}$. More precisely, let $R=\mathcal{H}|_{S}$ and $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$. These serve merely as representations for the functions in $R$ and variables in $L$, not their true definitions (which will be established shortly). Note now that by suppressing the unlabeled datapoints of $S$, we can equivalently represent elements of $R$ as sequences in $\mathcal{Y}^{n}$ and elements of $L$ as sequences in $(\mathcal{Y}\cup\{?\})^{n}$. In this view, each element of $L$ is precisely an element of $R$ which had exactly one entry replaced with a ``?"``?"``?"` ` ? ".
Now, to model transductive learning, fix an element $r\in R$ represented by $(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$. Then we will define $r$ to be a function depending upon the variables $\ell_{1},\ldots,\ell_{n}\in L$, where $\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$. Given assignments for each of the variables $\ell_{1},\ldots,\ell_{n}$ as values in $\mathcal{Y}$ — semantically, completions of their “?” entries — the node $r$ then outputs the value $\frac{1}{n}\cdot\sum_{i=1}^{n}d(y_{i},\ell_{i})$. The two crucial observations are as follows: an assignment of each $\ell\in L$ corresponds precisely to the action of a learner responding to a query at test time, and the output of node $r$ equals the error of a learner when $r$ is the ground truth.
@@
Theorem 3.6 establishes an exact compactness in learning with respect to a flexible class of metric loss functions. One may note, however, that some non-metric losses are of central importance to machine learning, including the squared error on compact subsets of $\mathbb{R}$ (which violates the triangle inequality) and the cross-entropy loss for finite-dimensional distributions (which is not symmetric). We now provide a modified form of Theorem 3.6 which captures these loss functions, in which the loss function $\ell_{\mathcal{Y}}$ is permitted to differ from the underlying metric $d$ on $\mathcal{Y}$. (E.g., such that $d$ is the usual Euclidean norm on a compact subset $\mathcal{Y}$ of $\mathbb{R}^{d}$, and $\ell_{\mathcal{Y}}$ is any continuous loss function.)
@@
Let $\mathcal{X}$ be an arbitrary domain, $(\mathcal{Y},d)$ a compact metric space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Let $\ell_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ be a loss function employed for learning that is continuous with respect to the metric $d$. Then the following are equivalent for any $m\colon\mathbb{R}_{>0}\to\mathbb{N}$:
  • 1.
    $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m$.
  • 2.
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the realizable case with transductive sample function $m$.
Proof.
We adopt precisely the perspective of Theorem 3.6, seeing transductive learning modeled as a variable assignment problem with the same variables $L$ and functions $R$. To invoke Theorem 3.3, it remains only to show that the functions $r\in R$ are proper. First note a continuous function from a compact space to $\mathbb{R}$ is automatically proper, as closed subsets of compact sets are compact. Now recall that each $r\in R$ is a sum of scaled copies of $\ell_{\mathcal{Y}}$ with one input fixed. As each such function is continuous, $r$ itself is continuous and thus proper. ∎
3.2 Realizable Learning With Improper Loss Functions
It is natural to ask whether the requirement that $\mathcal{Y}$ be a proper metric space is essential to Theorem 3.6 or merely an artifact of the proof. We now demonstrate the former: for arbitrary metric losses, the error rate of learning $\mathcal{H}$ can exceed that of all its finite projections by a factor of 2. Recall that $\xi_{\mathcal{H}}:\mathbb{N}\to\mathbb{R}_{\geq 0}$ denotes the transductive error rate of learning a class $\mathcal{H}$, i.e., $\xi_{\mathcal{H}}(n)$ denotes the error incurred by an optimal learner for $\mathcal{H}$ on (worst-case) samples of size $n$.